OpenStudy (anonymous):

can someone please help me on this question Glenn invested $18,000 at 5% interest compounded annually. How much interest will Glenn earn in 3 years? A. $20,837.25 B. $19,826.24 C. $2837.25 D. $2340.23

OpenStudy (amilapsn):

\[18000\times(\frac{ 5 }{ 100})^{3}\]

OpenStudy (anonymous):

i need help i cant do it

OpenStudy (anonymous):

a

OpenStudy (anonymous):

thanks cram

OpenStudy (triciaal):

simple interest, I = Principal,p * rate as a %, r/100* time ,t in years I = P RT/100 p = 18000; r = 5% = 5/100; t = 3 years plug the numbers compound interest I =P (1 + r)^t - P where r is the rate as a % and t = number of compounding periods plug the numbers and match the answer

OpenStudy (triciaal):

compound interest answer c @cram that would be the accumulated amount after the 3 years

OpenStudy (triciaal):

@sga the question asked for the interest